Abdullah saves $600 for new car. Currently Abdullahhave $(n+1). If his per week saving is $(n). How long he waits to save at least $700 to cover his tax’s.Also, Give the geometrical representation of solution. Where n is your arid number (for example if your arid number is 19-arid-1234 then choose n=1234). (2 points)
1
Expert's answer
2020-12-17T18:19:13-0500

If we consider t number of weeks which were required to collect 700 dollors, then:

"t\times n+(n+1)"=700, n is the weekly cost that it should be collected and n+1 is already collected money. If we find t:

"t=\frac{699}{n}-1" --> this is the number weeks that is required to collect 700 dollors.